# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 6 Flatness of the Hitchin map
We choose a trivialisation $X_{r}=\cup_{i\in I}V_{i}$ of the vector bundle $\mathcal{E}$ and the line bundle $\omega_{X_{r}}$. Then a first-order infinitesimal deformation (as a Higgs bundle) of $(E,\phi)$ can be described as a pair $(s_{ij},t_{i})$, where $s_{ij}\in\Gamma(V_{ij},\mathcal{E}ndE)$ and $t_{i}\in\Gamma(V_{i},\mathcal{E}ndE\otimes\omega_{X_{r}})$ satisfying $s_{ij}+s_{jk}=s_{ik}$ and $t_{i}-t_{j}=[s_{ij},\phi]$ (see [5, Theorem 2.3] and [6, Proposition 3.1.2]).
Let $\operatorname{Spec}(\mathbb{k}[\epsilon])\longrightarrow\mathcal{N}ilp^{sm,gen}_{0}$ be a map such that the image of the closed point is given by the Higgs bundle $(X_{r},E,\phi)$. Let us denote by the corresponding first-order infinitesimal deformation of the nilpotent Higgs bundle $(E,\phi)$ by $(E[\epsilon],\phi[\epsilon])$. We assume that $(E[\epsilon],\phi[\epsilon])$ is a nilpotent Higgs bundle. We define $E^{i}[\epsilon]:=\ker\leavevmode\nobreak\ (\phi[\epsilon])^{i}$ for $i\in[1,k+1]$. Since the induced flag $E^{\bullet}$ is flat over $\mathcal{N}ilp^{sm,gen}_{0}$ (see Lemma 6.7), therefore $E^{\bullet}[\epsilon]$ is flat over $\operatorname{Spec}(\mathbb{k}[\epsilon])$. Since $E^{i}[\epsilon]$ is flat over $\operatorname{Spec}(\mathbb{k}[\epsilon])$, therefore it is an extension of $E^{i}$ by $E^{i}$.
@@
Now it is straightforward to check that for $(s_{ij},t_{i})$ to be an infinitesimal deformation of $(E,\phi)$ as a nilpotent Higgs field, it has to satisfy the extra condition that $s_{ij}(E^{\bullet})\subseteq E^{\bullet-1}$, where $E^{\bullet}$ is the flag (6.2). This means that $s_{ij}\in\Gamma(V_{ij},SC(E,\phi))$ for all $i,j$ and therefore, the proposition follows. ∎
@@
  2. 2.
    The sub-stack $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic in $M^{cl}_{Gie}$ with respect to the relative 0-shifted log symplectic form which was defined in Theorem 4.9.
  3. 3.
@@
Proof of (1): To prove the surjectivity we use the spectral correspondence [1, Lemma 2.4]. Given any point $a_{\bullet}:=(a_{1},\dots,a_{n})\in B$, we consider the function $s(a_{\bullet}):=t^{n}+a_{1}t^{n-1}+\dots+a_{n-1}t+a_{n}$ on the total space $\operatorname{Tot}(\omega_{X_{0}})$. Here $t$ denotes the canonical section of $f^{*}\omega_{X_{0}}$, where $f:\operatorname{Tot}(\omega_{X_{0}})\longrightarrow X_{0}$ is the projection map. Then the vanishing locus $V(s(a_{\bullet}))$ defines a closed sub-scheme of $\operatorname{Tot}(\omega_{X_{0}})$. Notice that $\operatorname{Tot}(\omega_{X_{0}})$ is only quasi-projective and it is an open subscheme of $Z:=\mathbb{P}(\omega_{X_{0}}^{*}\oplus\mathcal{O}_{X_{0}})$. But since $s(a_{\bullet})$ is a monic polynomial, therefore the closure of $V(s(a_{\bullet}))$ in $Z$ is $V(s(a_{\bullet}))$ itself. In particular, $V(s(a_{\bullet}))$ is a closed subscheme in $Z$ such that $V(s(a_{\bullet}))\cap D_{\infty}=\emptyset$, where $D_{\infty}:=Z\setminus\operatorname{Tot}(\omega_{X_{0}})$. Therefore, by spectral correspondence, it follows that any rank $1$ locally free sheaf on $V(s(a_{\bullet}))$ corresponds to a Higgs bundle $(E,\phi)$ on the nodal curve $X_{0}$ whose characteristic polynomial is given by $a_{\bullet}:=(a_{1},\dots,a_{n})\in B$. Therefore, the Hitchin map is surjective.
@@
Proof of (2): We want to show that $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic over $S$. First, we notice that the smooth locus of the generic fibre of $\mathcal{N}ilp^{red}\longrightarrow S$ is isotropic, and its dimension is equal to $n^{2}(g-1)$, which is the same as the dimension of the stack of rank $n$ vector bundles on a smooth projective curve of genus $g$. Therefore, we have $\dim\mathcal{N}ilp^{red}_{0}\geq n^{2}(g-1)$. We want to show that the $\mathcal{N}ilp^{red,sm}_{0}$ is an isotropic substack in $M^{cl}_{Gie,0}$. The tangent complex at a point $(X_{r},E,\phi)$ of $\mathcal{N}ilp^{red,sm}_{0}$ is given by the derived global sections of (6.6). The tangent complex at the corresponding point of $M^{cl}_{Gie}$ is given by the derived global sections of the following complex (see 4.17 and Theorem 4.18).
$\mathcal{C}{(E,\phi)}:=[\mathcal{E}ndE\xrightarrow{[-,\phi]}\mathcal{E}ndE\otimes\omega_{X_{r}}]$ (6.8)
We want to show that the following composite morphism is homotopic to 00.
${R\Gamma(\mathcal{SC}(E,\phi))[1]}$${R\Gamma(\mathcal{C}(E,\phi))[1]}$${(R\Gamma(\mathcal{C}(E,\phi))[1])^{\vee}}$${(R\Gamma(\mathcal{SC}(E,\phi))[1])^{\vee}}$$\scriptstyle{\omega^{\flat}}$ (6.9)
@@
Notice that the morphism $\omega^{\flat}:R\Gamma(\mathcal{C}(E,\phi))[1]\to(R\Gamma(\mathcal{C}(E,\phi))[1])^{\vee}$ is given by the log-symplectic pairing (Theorem 4.9). If we show that the composite morphism 6.9 is homotopic to 00, this will imply that the $\mathcal{N}ilp^{red,sm}_{0}$ is an isotropic substack. To show this, we choose a cover $X_{r}=\cup_{i\in I}V_{i}$ for which the vector bundle $E$ and the line bundle $\omega_{X_{r}}$ trivialize. With respect to this cover, the complex $R\Gamma(\mathcal{SC}(E,\phi))[1]$ is quasi-isomorphic to the following Čech complex
@@
where the term $\prod SC(E,\phi)({V_{i}})$ is in degree $-1$.
Notice that the composite map can only be non-zero on three cohomologies: in the degree $-1$, 00 and $1$.
@@
is given by $\{(s_{ij},t_{i})\}_{i,j\in I}$, where $(s_{ij},t_{i})$, where $s_{ij}\in\Gamma(V_{ij},\mathcal{E}ndE)$ and $t_{i}\in\Gamma(V_{i},\mathcal{E}ndE\otimes\omega_{X_{r}})$ such that $s_{ij}(E^{\bullet})\subseteq E^{\bullet-1}$, where $E^{\bullet}$ is the flag (6.2). Then from (6.9) it follows that the pairing of two elements $\{(s_{ij},t_{i})\}$ and $\{(s^{\prime}_{ij},t^{\prime}_{i})\}$ is given by
@@
We notice that $V_{ij}$’s also contain at most one node. Suppose, $V_{ij}$ contains a node $p$. Then $V_{ij}=V^{1}_{ij}\coprod V^{2}_{ij}$, the union of two smooth irreducible components of $V_{ij}$. We denote by $V^{o}_{ij}:=V_{ij}\setminus p$, $V^{1,o}_{ij}:=V^{1}_{ij}\setminus p$ and $V^{2,o}_{ij}:=V^{2}_{ij}\setminus p$. We consider the restriction of the section $T_{ij}$ to the two open subsets $V^{1,o}_{ij}$ and $V^{2,o}_{ij}$. Notice that the restrictions of the flag (6.2) to these two open subsets are all sub-bundles. We notice that $SC(E|_{V^{1,o}_{ij}},\phi)$ is the nilpotent part of the parabolic sub-algebra of $\mathcal{E}nd(E|_{V^{1,o}_{ij}})$ given by the natural flag of sub-bundles induced by $\phi$. Therefore it is clear that the trace pairing is 00 i.e., $(T_{ij})|_{V^{1,o}_{ij}}=0$. Similarly, $(T_{ij})|_{V^{2,o}_{ij}}=0$. Since $V^{o}_{ij}$ is a dense open subset of $V_{ij}$, therefore, by continuity, $T_{ij}=0$.
@@
Case 2: the induced map $H^{-1}(\omega^{\flat})$: In this case, we look at the pairing of $\{t_{i}\in SC(E,\phi)\otimes\omega_{X_{r}}\}$ and $\{s_{ij}\in SC(E,\phi)\}$. The pairing is given by $\{\operatorname{tr}(t_{i}\circ s_{ij})\}$. By similar arguments as above, we can show that $\operatorname{tr}(t_{i}\circ s_{ij})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \forall i,j\in I$.
Case 3: the induced map $H^{1}(\omega^{\flat})$: In this case, we look at the pairing of $\{t_{ij}\in(SC(E,\phi)\otimes\omega_{X_{r}})(V_{ij})\}$ and $\{s_{j}\in SC(E,\phi)(V_{j})\}$. The pairing is given by $\{\operatorname{tr}(t_{ij}\circ s_{j})\}$. By similar arguments as above, we can show that $\operatorname{tr}(t_{ij}\circ s_{j})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \forall i,j\in I$.
@@
Proof of (3) From (2), it follows that $\dim\mathcal{N}ilp^{red}_{0}=n^{2}(g-1)$. Now from Lemma 8.11, it follows that the dimension of $M^{cl}_{Gie}$ is equal to $22n^{2}(g-1)+$ 2. Once we know the dimension, it is easy to see that $M^{cl}_{Gie}$ is a local complete intersection (Remark 8.13). Since $M^{cl}_{Gie}$ is a local complete intersection, and all the fibres of the Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ are of dimension equal to $n^{2}(g-1)$, therefore by the miraculous flatness criterion [13, Exercise 10.9, page 276], it follows that the Hitchin map is flat. ∎
7 On the relative logarithmic Dolbeault moduli over $\overline{\mathcal{M}_{g}}$
In this section, we will construct the derived moduli stack $M_{g}^{Dol}$ of Higgs bundles with coefficients in the relative dualizing sheaf along the fibers of the universal curve over the moduli stack $\mathcal{M}_{g}^{ss}$ of semistable curves of genus $g\geq 2$. We will show that there is a relative 00-shifted symplectic form on $M_{g}^{Dol}/\mathcal{M}_{g}^{ss}$ which can also be interpreted as a relative 00-shifted log-symplectic form on $M_{g}^{Dol}$, relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$. Let us first recall some notations.
@@ 7 On the relative logarithmic Dolbeault moduli over $\overline{\mathcal{M}_{g}}$
$T\mapsto\left\{\begin{array}[]{@{}ll@{}}\text{families of stable curves}\\ \text{ of genus}\leavevmode\nobreak\ g\leavevmode\nobreak\ \text{over}\leavevmode\nobreak\ T\end{array}\right\}$ (7.1)
@@
It is well known that if $g\geq 2$, then $\overline{\mathcal{M}_{g}}$ is a smooth and proper Deligne-Mumford stack. The locus of nodal curves forms a normal crossing divisor in $\overline{\mathcal{M}_{g}}$ and so gives a divisorial log structure on $\overline{\mathcal{M}_{g}}$. On the other hand, by a well known result of S. Mochizuki [23, §3B.], given any family of semi-stable curves $\mathcal{C}\to T$, there is a natural log structure on $\mathcal{C}$ and also on the base scheme $T$ such that the projection map is a log-smooth map and the relative logarithmic cotangent bundle is the relative dualizing sheaf of $\mathcal{C}/T$. We call these log structures the basic log structure package on $\mathcal{C}/T$. It is known that the log structure on $\overline{\mathcal{M}_{g}}$ which is part of the basic log structure package for the universal curve $\mathcal{C}_{g}$ over $\overline{\mathcal{M}_{g}}$ agrees with the divisorial log-structure on $\overline{\mathcal{M}_{g}}$ (see the comment after Lemma 4.4 in [21] and the proof of our Proposition 7.3).
@@
$T\mapsto\left\{\begin{array}[]{@{}ll@{}}\text{families of semi-stable curves}\\ \text{of genus}\leavevmode\nobreak\ g\leavevmode\nobreak\ \text{over}\leavevmode\nobreak\ T\end{array}\right\}$ (7.2)
@@
It is well known that if $g\geq 2$, ${\mathcal{M}^{ss}_{g}}$ is a smooth Artin stack [39, 0E72, Lemma 21.5] of finite type.
7.1 Log structures on ${\mathcal{M}^{ss}_{g}}$
The locus of singular curves in ${\mathcal{M}^{ss}_{g}}$ again forms a normal crossing divisor, and hence induces a log structure on ${\mathcal{M}^{ss}_{g}}$. We denote this divisor by $\partial\mathcal{M}^{ss}_{g}$. Let us denote the universal curve over ${\mathcal{M}^{ss}_{g}}$ by $\mathcal{D}_{g}$. There is the “stabilization morphism” $\pi:\mathcal{M}^{ss}_{g}\longrightarrow\overline{\mathcal{M}_{g}}$ such that the induced morphism $\mathcal{D}_{g}\longrightarrow\pi^{*}\mathcal{C}_{g}$ is the universal modification morphism over $\mathcal{M}^{ss}_{g}$.
7.1.1 Versal deformations of the stabalization morphism

### v3

@@ 6. Flatness of the Hitchin map
We choose a trivialisation $X_{r}=\cup_{i\in I}V_{i}$ of the vector bundle $\mathcal{E}$ and the line bundle $\omega_{X_{r}}$. Then a first-order infinitesimal deformation (as a Higgs bundle) of $(E,\phi)$ can be described as a pair $(s_{ij},t_{i})$, where $s_{ij}\in\Gamma(V_{ij},\mathcal{E}\mathsf{nd}E)$ and $t_{i}\in\Gamma(V_{i},\mathcal{E}\mathsf{nd}E\otimes\omega_{X_{r}})$ satisfying $s_{ij}+s_{jk}=s_{ik}$ and $t_{i}-t_{j}=[s_{ij},\phi]$ (see [6, Theorem 2.3] and [7, Proposition 3.1.2]).
Let $\operatorname{Spec}(\mathbb{k}[\epsilon])\longrightarrow\mathcal{N}ilp^{sm,gen}_{0}$ be a map such that the image of the closed point is given by the Higgs bundle $(X_{r},E,\phi)$. Let us denote by the corresponding first-order infinitesimal deformation of the nilpotent Higgs bundle $(E,\phi)$ by $(E[\epsilon],\phi[\epsilon])$. We assume that $(E[\epsilon],\phi[\epsilon])$ is a nilpotent Higgs bundle. We define $E^{i}[\epsilon]:=\ker\penalty 10000\ (\phi[\epsilon])^{i}$ for $i\in[1,k+1]$. Since the induced flag $E^{\bullet}$ is flat over $\mathcal{N}ilp^{sm,gen}_{0}$ (see Lemma 6.7), therefore $E^{\bullet}[\epsilon]$ is flat over $\operatorname{Spec}(\mathbb{k}[\epsilon])$. Since $E^{i}[\epsilon]$ is flat over $\operatorname{Spec}(\mathbb{k}[\epsilon])$, therefore it is an extension of $E^{i}$ by $E^{i}$.
@@
Now it is straightforward to check that for $(s_{ij},t_{i})$ to be an infinitesimal deformation of $(E,\phi)$ as a nilpotent Higgs field, it has to satisfy the extra condition that $s_{ij}(E^{\bullet})\subseteq E^{\bullet-1}$ and $t_{i}(E^{\bullet})\subseteq E^{\bullet-1}\otimes\omega_{X_{r}}$, where $E^{\bullet}$ is the flag (6.2). This means that $s_{ij}\in\Gamma(V_{ij},SC(E,\phi))$ for all $i,j$ and therefore, the proposition follows. ∎
@@
  2. (2)
    The sub-stack $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic in $M^{cl}_{Gie}$ with respect to the relative 0-shifted log symplectic form defined in Theorem 4.11.
  3. (3)
@@
Proof of (1): Let $a_{\bullet}:=(a_{1},\dots,a_{n})\in B$ be any point. This point $a_{\bullet}$ can either lie over the closed fibre of the map $B\to S$ or outside the closed fibre. If it lies outside the closed fibre, then it lies in the image of the Hitchin map. This fact follows from the spectral correspondence ([37, Theorem 6.11]). If $a_{\bullet}$ lies over the closed fibre, we again use a version of spectral correspondence for the Hitchin map for nodal curves [2, Lemma 2.4]. Let us consider the function $s(a_{\bullet}):=t^{n}+a_{1}t^{n-1}+\dots+a_{n-1}t+a_{n}$ on the total space $\operatorname{Tot}(\omega_{X_{0}})$. Here $t$ denotes the canonical section of $f^{*}\omega_{X_{0}}$, where $f:\operatorname{Tot}(\omega_{X_{0}})\longrightarrow X_{0}$ is the projection map. Then the vanishing locus $V(s(a_{\bullet}))$ defines a closed sub-scheme of $\operatorname{Tot}(\omega_{X_{0}})$. Notice that $\operatorname{Tot}(\omega_{X_{0}})$ is only quasi-projective and it is an open subscheme of $Z:=\mathbb{P}(\omega_{X_{0}}^{*}\oplus\mathcal{O}_{X_{0}})$. But since $s(a_{\bullet})$ is a monic polynomial, therefore the closure of $V(s(a_{\bullet}))$ in $Z$ is $V(s(a_{\bullet}))$ itself. In particular, $V(s(a_{\bullet}))$ is a closed subscheme in $Z$ such that $V(s(a_{\bullet}))\cap D_{\infty}=\emptyset$, where $D_{\infty}:=Z\setminus\operatorname{Tot}(\omega_{X_{0}})$. Therefore, by spectral correspondence, it follows that any rank $1$ locally free sheaf on $V(s(a_{\bullet}))$ corresponds to a Higgs bundle $(E,\phi)$ on the nodal curve $X_{0}$ whose characteristic polynomial is given by $a_{\bullet}:=(a_{1},\dots,a_{n})\in B$. Therefore, the Hitchin map is surjective.
@@
Proof of (2): We want to show that $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic over $S$. First, we notice that the smooth locus of the generic fibre of $\mathcal{N}ilp^{red}\longrightarrow S$ is isotropic, and its dimension is equal to $n^{2}(g-1)$, which is the same as the dimension of the stack of rank $n$ vector bundles on a smooth projective curve of genus $g$. Therefore, we have $\dim\mathcal{N}ilp^{red}_{0}\geq n^{2}(g-1)$. We want to show that the $\mathcal{N}ilp^{red,sm}_{0}$ is an isotropic substack in $M^{cl}_{Gie,0}$. The tangent complex at a point $(X_{r},E,\phi)$ of $\mathcal{N}ilp^{red,sm}_{0}$ is given by the derived global sections of (6.6). The tangent complex at the corresponding point of $M^{cl}_{Gie}$ is given by the derived global sections of the following complex (see (4.17) and Theorem 4.18).
(6.8) $\mathcal{C}{(E,\phi)}:=[\mathcal{E}\mathsf{nd}E\xrightarrow{[-,\phi]}\mathcal{E}\mathsf{nd}E\otimes\omega_{X_{r}}]$
We want to show that the following composite morphism is homotopic to 0.
(6.9) $R\Gamma(\mathcal{SC}(E,\phi))[1]$$R\Gamma(\mathcal{C}(E,\phi))[1]$$(R\Gamma(\mathcal{C}(E,\phi))[1])^{\vee}$$(R\Gamma(\mathcal{SC}(E,\phi))[1])^{\vee}$$\scriptstyle{\omega^{\flat}}$
@@
Notice that the morphism $\omega^{\flat}:R\Gamma(\mathcal{C}(E,\phi))[1]\to(R\Gamma(\mathcal{C}(E,\phi))[1])^{\vee}$ is given by the logarithmic-symplectic pairing (Theorem 4.11). If we show that the composite morphism (6.9) is homotopic to 0, this will imply that $\mathcal{N}ilp^{red,sm}_{0}$ is an isotropic substack. To show this, we choose a cover $X_{r}=\underset{i\in I}{\cup}V_{i}$ for which the vector bundle $E$ and the line bundle $\omega_{X_{r}}$ trivialize. With respect to this cover, the complex $R\Gamma(\mathcal{SC}(E,\phi))[1]$ is quasi-isomorphic to the following Čech complex
@@
where the term $\prod SC(E,\phi)({V_{i}})$ is in degree 1-1.
Notice that the composite map can only be non-zero on three cohomologies: in the degree 1-1, 0 and $1$.
@@
is given by $\{(s_{ij},t_{i})\}_{i,j\in I}$, where $(s_{ij},t_{i})$, where $s_{ij}\in\Gamma(V_{ij},\mathcal{E}\mathsf{nd}E)$ and $t_{i}\in\Gamma(V_{i},\mathcal{E}\mathsf{nd}E\otimes\omega_{X_{r}})$ such that $s_{ij}(E^{\bullet})\subseteq E^{\bullet-1}$, and $t_{i}(E^{\bullet})\subseteq E^{\bullet-1}\otimes\omega_{X_{r}}$, where $E^{\bullet}$ is the flag (6.2). Then from (6.9) it follows that the pairing of two such elements $\{(s_{ij},t_{i})\}$ and $\{(s^{\prime}_{ij},t^{\prime}_{i})\}$ is given by
@@
We notice that $V_{ij}$’s also contain at most one node. Suppose, $V_{ij}$ contains a node $p$. Then $V_{ij}=V^{1}_{ij}\coprod V^{2}_{ij}$, the union of two smooth irreducible components of $V_{ij}$. We denote by $V^{o}_{ij}:=V_{ij}\setminus p$, $V^{1,o}_{ij}:=V^{1}_{ij}\setminus p$ and $V^{2,o}_{ij}:=V^{2}_{ij}\setminus p$. We consider the restriction of the section $T_{ij}$ to the two open subsets $V^{1,o}_{ij}$ and $V^{2,o}_{ij}$. Notice that the restrictions of the flag (6.2) to these two open subsets are all sub-bundles. We notice that $SC(E|_{V^{1,o}_{ij}},\phi)$ is the nilpotent part of the parabolic sub-algebra of $\mathcal{E}\mathsf{nd}(E|_{V^{1,o}_{ij}})$ given by the natural flag of sub-bundles induced by $\phi$. Therefore it is clear that the trace pairing is 0 i.e., $(T_{ij})|_{V^{1,o}_{ij}}=0$. Similarly, $(T_{ij})|_{V^{2,o}_{ij}}=0$. Since $V^{o}_{ij}$ is a dense open subset of $V_{ij}$, therefore, by continuity, $T_{ij}=0$.
@@
Case 2: the induced map $H^{-1}(\omega^{\flat})$: In this case, we look at the pairing of $\{t_{i}\in SC(E,\phi)\otimes\omega_{X_{r}}\}$ and $\{s_{ij}\in SC(E,\phi)\}$. The pairing is given by $\{\operatorname{tr}(t_{i}\circ s_{ij})\}$. By similar arguments as above, we can show that $\operatorname{tr}(t_{i}\circ s_{ij})=0\penalty 10000\ \penalty 10000\ \forall i,j\in I$.
Case 3: the induced map $H^{1}(\omega^{\flat})$: In this case, we look at the pairing of $\{t_{ij}\in(SC(E,\phi)\otimes\omega_{X_{r}})(V_{ij})\}$ and $\{s_{j}\in SC(E,\phi)(V_{j})\}$. The pairing is given by $\{\operatorname{tr}(t_{ij}\circ s_{j})\}$. By similar arguments as above, we can show that $\operatorname{tr}(t_{ij}\circ s_{j})=0\penalty 10000\ \penalty 10000\ \forall i,j\in I$.
@@
Proof of (3) From (2), it follows that $\dim\mathcal{N}ilp^{red}_{0}=n^{2}(g-1)$. Now from Lemma A.12, it follows that the dimension of $M^{cl}_{Gie}$ is equal to $2n^{2}(g-1)+2$. Once we know the dimension, it is easy to see that $M^{cl}_{Gie}$ is a local complete intersection (Corollary A.13). Since $M^{cl}_{Gie}$ is a local complete intersection, and all the fibres of the Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ are of dimension equal to $n^{2}(g-1)$, therefore by the miraculous flatness criterion [13, Exercise 10.9, page 276], it follows that the Hitchin map is flat. ∎
7. On the relative logarithmic Dolbeault moduli over $\overline{\mathcal{M}_{g}}$
In this section, we will construct the derived moduli stack $M_{g}^{Dol}$ of Higgs bundles with coefficients in the relative dualizing sheaf along the fibers of the universal curve over the moduli stack $\mathcal{M}_{g}^{ss}$ of semistable curves of genus $g\geq 2$. We will show that there is a relative 0-shifted symplectic form on $M_{g}^{Dol}/\mathcal{M}_{g}^{ss}$ which can also be interpreted as a relative 0-shifted log-symplectic form on $M_{g}^{Dol}$, relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$. Let us first recall some notations.
@@ 7. On the relative logarithmic Dolbeault moduli over $\overline{\mathcal{M}_{g}}$
(7.1) $T\mapsto\left\{\begin{array}[]{@{}ll@{}}\text{families of stable curves}\\ \text{ of genus}\penalty 10000\ g\penalty 10000\ \text{over}\penalty 10000\ T\end{array}\right\}$
@@
It is well known that if $g\geq 2$, then $\overline{\mathcal{M}_{g}}$ is a smooth and proper Deligne-Mumford stack. The locus of nodal curves forms a normal crossing divisor in $\overline{\mathcal{M}_{g}}$ and so gives a divisorial log structure on $\overline{\mathcal{M}_{g}}$. On the other hand, by a well known result of S. Mochizuki [24, §3B.], given any family of semi-stable curves $\mathcal{C}\to T$, there is a natural log structure on $\mathcal{C}$ and also on the base scheme $T$ such that the projection map is a log-smooth map and the relative logarithmic cotangent bundle is the relative dualizing sheaf of $\mathcal{C}/T$. We call these log structures the basic log structure package on $\mathcal{C}/T$. It is known that the log structure on $\overline{\mathcal{M}_{g}}$ which is part of the basic log structure package for the universal curve $\mathcal{C}_{g}$ over $\overline{\mathcal{M}_{g}}$ agrees with the divisorial log-structure on $\overline{\mathcal{M}_{g}}$ (see the comment after Lemma 4.4 in [15] and the proof of our Proposition 7.3).
@@
(7.2) $T\mapsto\left\{\begin{array}[]{@{}ll@{}}\text{families of semi-stable curves}\\ \text{of genus}\penalty 10000\ g\penalty 10000\ \text{over}\penalty 10000\ T\end{array}\right\}$
@@
It is well known that if $g\geq 2$, $\mathcal{M}^{ss}_{g}$} is a smooth Artin stack [41, 0E72, Lemma 21.5] of finite type.
7.1. Log structures on $\mathcal{M}^{ss}_{g}$}
The locus of singular curves in $\mathcal{M}^{ss}_{g}$} again forms a normal crossing divisor, and hence induces a log structure on $\mathcal{M}^{ss}_{g}$}. We denote this divisor by $\partial\mathcal{M}^{ss}_{g}$. Let us denote the universal curve over $\mathcal{M}^{ss}_{g}$} by $\mathcal{D}_{g}$. There is the “stabilization morphism” $\pi:\mathcal{M}^{ss}_{g}\longrightarrow\overline{\mathcal{M}_{g}}$ such that the induced morphism $\mathcal{D}_{g}\longrightarrow\pi^{*}\mathcal{C}_{g}$ is the universal modification morphism over $\mathcal{M}^{ss}_{g}$. The existence of the universal modification follows from a well-known result called the stable reduction theorem [41, Lemma 109.23.4.]. The existence of the stabilisation morphism follows from the universal property of $\overline{\mathcal{M}_{g}}$ and the universal modification morphism.
7.1.1. Versal deformations of the stabalization morphism